# A Survey of Cognitive Distortion Detection and Classification in NLP

### v1

@@ Abstract
As interest grows in the application of natural language processing techniques to mental health, a growing body of work explores the automatic detection and classification of cognitive distortions. cognitive distortions are habitual patterns of negatively biased or flawed thinking that distort how people perceive events, judge themselves, and react to the world around them. Identifying and addressing them is an important part of therapy. Despite its momentum, the field remains fragmented, with inconsistencies in cognitive distortion taxonomies, task formulations, and evaluation practices. This survey reviews 38 studies spanning two decades, providing a structured overview of datasets, modelling approaches, and evaluation strategies. We provide a consolidated cognitive distortion taxonomy reference, summarise common task setups, and highlight open challenges to support more coherent and reproducible research in this emerging area.
@@ 1 Introduction
Common examples include Catastrophising (‘I let them down, so they’ll never trust me again’), Mind Reading (‘They haven’t replied, so they must be angry at me’), and All or Nothing Thinking (‘If I don’t get this right the first time, I’m a complete failure’). [^1] A complete list of cognitive distortion definitions, examples, and synonyms is provided in Table 3, with additional psychological background in Appendix B. These patterns often appear intuitive or harmless at first glance, but they have been shown to maintain and exacerbate conditions like depression, anxiety, and post-traumatic stress disorder. In therapeutic settings, recognising and reframing such distortions is a core goal of cognitive behavioural therapy. Crucially, these distortions are primarily expressed through language, making them well-suited to computational modelling. Recent work in natural language processing has begun to explore the automatic detection and classification of cognitive distortions, with applications ranging from clinical decision support tools to mental health chatbots, journaling tools, and triage systems. Studies have shown that incorporating cognitive distortion-level features can improve outcomes in related tasks such as depression detection (Wang et al., 2023b), complementing more traditional sentiment or topic based approaches. By identifying distorted cognitive patterns in everyday text, natural language processing systems may support more timely, personalised, and psychologically-informed interventions.
@@
To our knowledge, the most comparable prior survey is by Suputra et al. (2023), which examined 12 studies and provided an initial synthesis of modelling approaches used at the time. Our work expands on this by covering 38 publications, including recent preprints, and offering a structured overview of the computational landscape for cognitive distortion detection and classification. This survey aims to provide a clear and practical entry point for researchers engaged in the growing field of cognitive distortion detection and classification from a computational perspective. Our focus is firmly on how computational methods approach these tasks, without seeking to redefine or adjudicate psychological constructs themselves. The contributions in this paper are threefold. (i) It provides a consolidated reference of the cognitive distortion taxonomies used across computational studies, highlighting inconsistencies and listing common synonyms (Tables 1 and 3). (ii) It defines and analyses task setups, datasets, computational methods, and approaches to performance evaluation, highlighting key patterns and gaps (§2, §4, §5, §6). (iii) It identifies open challenges and proposes concrete directions to guide future research (§7). In doing so, this paper aims to enable more consistent, comparable, and reproducible work in this emerging field.
@@ 3 Taxonomies of cognitive distortions in natural language processing
Despite the shared objective of classifying cognitive distortions, studies seeking to label cognitive distortions have adopted diverse taxonomies. Without standardisation, it is hard to keep annotations consistent, compare models properly, or clearly interpret the results. Table 1 illustrates this fragmentation by mapping which cognitive distortion categories are recognised across the papers surveyed. While some categories, such as Labelling and Should Statements, are commonly used, others are inconsistently applied or redefined in various ways. Early conceptions of cognitive distortions were proposed by Beck (1963), who described patterns of dysfunctional thinking with examples such as arbitrary inference and overgeneralisation. This work was later popularised by Burns (1999), whose ten-category taxonomy is frequently cited in psychology literature [^2] While Burns (1999) is not the only taxonomy of cognitive distortions used in psychological contexts, and clinical practice often employs broader or alternative frameworks, our focus is on the inconsistency in how these definitions are interpreted and applied within computational research.. However, computational studies do not uniformly follow this framework.
@@ 4.3 Annotation Strategies
Annotation strategies vary considerably across datasets. Clinical and literature-derived corpora typically rely on expert annotators, prioritising label quality at the expense of scalability. In contrast, social media and crowd-sourced datasets often involve non-expert annotators, sometimes supported by brief, domain-specific training from qualified psychologists to improve consistency. While these strategies enable large-scale annotation, reliably classifying cognitive distortions remains challenging. The task demands subtle, often subjective judgements, and studies consistently report low inter-annotator agreement, particularly when annotators lack deeper domain expertise. To mitigate this, some works adopt strict inclusion criteria, only retaining examples where annotators fully agree on the label or a subset of distortion types (Aureus et al., 2021; Shickel et al., 2020). Though this approach improves label precision, it risks introducing bias by systematically excluding ambiguous or borderline cases - which are arguably the most reflective of real-world cognitive distortion occurrences. Balancing scalability, annotation quality, and representativeness therefore remains a core challenge in cognitive distortion dataset construction. Moreover, these annotation choices have important implications for how inter-annotator agreement is evaluated - a point we revisit (§6).
@@ 5 Modelling Approaches
This section outlines computational approaches to cognitive distortion detection and classification, grouped into six methodological categories that reflect major developments in the field. Our overview is informed by experimental findings from the surveyed literature, detailed in the supplementary materials.
@@ 5.2 Traditional Machine Learning (Feature-based)
As more annotated corpora became available, early rule-based systems gave way to feature-based statistical models. These approaches combined classic classifiers, such as logistic regression (LR) and support vector machines (SVMs), with engineered features such as Linguistic Inquiry and Word Count (LIWC) scores and Term Frequency-Inverse Document Frequency (TF-IDF) vectors. Simms et al. (2017) demonstrated that LR trained on LIWC features performed well on the detection task using Tumblr data, while Shickel et al. (2020) found that TF–IDF with LR outperformed CNNs on a 15-class single-label synthetic dataset (F1 = 0.68), indicating that shallow linguistic cues can remain competitive even in classification tasks. Similarly, Shreevastava and Foltz (2021) used SVM with smooth inverse frequency (SIF) embeddings, achieving strong performance (F1 = 0.77) on the detection task despite the insensitivity of SIF to word order.
@@ 5.3 Deep Learning with Static Embeddings
Static word embeddings such as Word2Vec (Mikolov et al., 2013) and GloVe (Pennington et al., 2014) introduced vector-based representations of words derived from co-occurrence patterns. These effectively captured word similarity but failed to account for contextual nuance (e.g. ‘riverbank’ vs ‘bank’). Despite this limitation, such embeddings formed the backbone of early deep-learning based cognitive distortion classifiers. For example, Rojas-Barahona et al. (2018) combined GloVe with CNNs for multi-label cognitive distortion classification, outperforming traditional models. Similarly, Mostafa et al. (2021) trained LSTM models using GloVe vectors to classify different cognitive distortion types, although their findings highlighted concerns around overfitting due to the use of limited and synthetic data. To address data sparsity in Arabic texts, Alhaj et al. (2022) applied contextual topic modelling (CTM), integrating static embeddings with domain-specific topic information via BERTopic - an approach that proved helpful in low-resource settings. While static embeddings offered richer representations than earlier methods, their inability to capture context, especially for polysemous words and subtle pragmatic distinctions, ultimately led to the rise of contextual models.
@@ 5.4 Transformer-based Architectures
Transformers, particularly BERT (Bidirectional Encoder Representations from Transformers) (Devlin et al., 2019), introduced contextual embeddings that capture word meaning based on the surrounding sentence, marking a significant leap forward in many natural language processing tasks, including cognitive distortion detection and classification. Shreevastava and Foltz (2021) used fine-tuned Sentence-BERT (SBERT) for binary cognitive distortion detection, showing notable improvements over earlier models. Similarly, Lybarger et al. (2022) demonstrated that incorporating conversational history further enhances classification performance on therapy dialogues. Domain-adapted transformers like MentalBERT (Ji et al., 2022), which are pre-trained on mental-health-related data, showed benefits over general-purpose models, while augmentation techniques such as mixup (Zhang et al., 2018) were explored to improve performance on rare cognitive distortion classes (minor improvements). Maddela et al. (2023) found that models like RoBERTa performed better than larger language models such as GPT-3.5 for this task, though distinguishing between closely related distortions remained a challenge. Overall, transformers brought significant gains in robustness and accuracy, but these came at the cost of reduced interpretability and a higher risk of overfitting, especially in datasets with extreme class imbalance. These limitations set the stage for the emergence of prompt-based models.
@@ 6 Evaluation
Despite progress in cognitive distortion classification, evaluation practices remain inconsistent, with studies differing markedly in their choice and reporting of metrics, which hinders comparability. Although F1 score is the most commonly used metric, distinctions between macro, micro, and weighted variants are frequently overlooked - a significant issue for class-imbalanced datasets where per-class performance is critical. Some studies now report AUPRC (Area Under the Precision–Recall Curve) to account for skewed label distributions, providing a more informative measure of performance on rare cognitive distortion types (Ding et al., 2022). Nonetheless, per-class metrics are still underreported, hiding weaknesses in addressing infrequent distortions. Similarly, dataset quality is often inconsistently assessed. Inter-annotator agreement (IAA) is either inconsistently measured or reported using incomparable metrics. While Cohen’s Kappa, which adjusts for chance, is typically more appropriate, many studies instead rely on raw agreement or pairwise F1, blurring the line between annotation reliability and model performance. Baseline comparisons are further complicated by inconsistent cognitive distortion taxonomies and datasets. We therefore recommend that researchers report per-class F1 scores, provide transparent IAA using chance-adjusted measures, and apply greater scrutiny to dataset origins.
@@ 7.1 Inconsistent cognitive distortion Taxonomies
A longstanding challenge is the lack of a standardised taxonomy for cognitive distortions. While foundational frameworks such as the Burns ten-category list (Burns, 1999) are commonly cited, computational studies diverge significantly in how they define, split, or rename distortion types. For instance, Jumping to Conclusions is frequently subdivided into Mind Reading and Fortune Telling, and terms such as All or Nothing Thinking appear under multiple aliases (e.g., Black and White Thinking, Polarised Thinking), making it difficult to compare models across studies, reproduce results, or interpret outputs reliably. These inconsistencies also affect annotation quality, as ambiguous or overly granular label sets introduce subjectivity and reduce inter-annotator agreement - an issue compounded by the lack of formal guidance on taxonomy use. While we provide a consolidated reference (Table 1), we do not claim it to be definitive. In the meantime, we encourage researchers to (i) report the source and rationale for their chosen taxonomy, (ii) avoid expanding taxonomies without empirical justification, and (iii) prioritise those grounded in clinical consensus where possible - with Burns’ taxonomy as a sensible default unless there is a well-founded reason to depart from it.
@@ 7.2 Data Scarcity, Imbalance, and Annotation Limits
The field remains constrained by a lack of large, high-quality datasets that capture authentic, context-rich examples of distorted thinking. Clinical corpora are scarce and often inaccessible due to privacy constraints, while many widely used datasets are synthetic, crowd-sourced, or compiled from multiple domains, and may lack the nuance and ambiguity of real-world language. This limits the complexity of distortions that models can learn and typically results in heavy class imbalance, with rare distortion types being underrepresented or excluded altogether. Although augmentation strategies such as mixup or back-translation offer minor gains for rare categories (Ding et al., 2022), a deeper issue lies in the ceiling imposed by annotation reliability itself. In some settings, even expert annotators show limited agreement, particularly for subtle or overlapping categories. For instance, Tauscher et al. (2023) report that for the presence of ‘Any Distortion’ in a text, human–human F1 agreement was 0.63, while a fine-tuned BERT model reached 0.62 - a very small difference. This suggests that for certain formulations, such as binary detection or high-frequency classes, current models may already be approaching the upper bound set by annotation quality. It also reinforces the need for clearer task definitions and more consistent annotation protocols before investing in model complexity. We recommend that future work (i) continues to report human–model comparison metrics where possible, (ii) prioritises the release of multilingual and multi-domain corpora, and (iii) explicitly documents how annotation disagreements are resolved. Without such efforts, further gains in model performance may reflect noise-fitting rather than genuine progress.
@@ 7.3 Overreliance on Short Text
The vast majority of existing datasets frame cognitive distortion detection at the sentence or single-utterance level. This simplifies annotation and model design but introduces strong limitations, as many distortions are context-dependent or only weakly signalled lexically. By stripping away discourse-level cues, models are forced to rely on surface-level patterns and may perform poorly on more ambiguous cases. Predictably, this has contributed to the dominance of distortion types with overt markers (e.g., Should Statements) in both datasets and model outputs. Empirical studies confirm the value of richer context: Lybarger et al. (2022) showed that including prior conversational turns improved detection F1 from 0.68 to 0.73, while frameworks like ERD (Lim et al., 2024) achieve greater interpretability by explicitly reasoning over multi-sentence inputs. Still, most benchmarks continue to prioritise short-form inputs. Moving forward, we argue that context-aware models should become the norm rather than the exception. In parallel, new datasets should prioritise multi-turn conversations, real patient narratives, and longer-form content that more closely mirrors therapeutic language.

### v2

@@ Abstract
As interest grows in applying natural language processing techniques to mental health, an expanding body of work explores the automatic detection and classification of cognitive distortions. cognitive distortions are habitual patterns of negatively biased or flawed thinking that distort how people perceive events, judge themselves, and react to the world. Identifying and addressing them is a central goal of therapy. Despite this momentum, the field remains fragmented, with inconsistencies in cognitive distortion taxonomies, task formulations, and evaluation practices limiting comparability across studies. This survey presents the first comprehensive review of 38 studies spanning two decades, mapping how cognitive distortions have been implemented in computational research and evaluating the methods applied. We provide a consolidated cognitive distortion taxonomy reference, summarise common task setups, and highlight persistent challenges to support more coherent and reproducible research. Alongside our review, we introduce practical resources, including curated evaluation metrics from surveyed papers, a standardised datasheet template, and an ethics flowchart, available online. [^1] https://github.com/archiesage/cognitive-distortion-nlp-survey
@@ 1 Introduction
Common examples [^2] A complete list of cognitive distortion definitions, examples, and synonyms is provided in Table 3, with additional psychological background in Appendix B. include Catastrophising (‘I let them down, so they’ll never trust me again’), Mind Reading (‘They haven’t replied, so they must be angry at me’), and All or Nothing Thinking (‘If I don’t get this right the first time, I’m a complete failure’). These patterns often appear intuitive or harmless at first, but they have been shown to maintain and exacerbate conditions like depression, anxiety, and post-traumatic stress disorder. In therapeutic settings, recognising and reframing such distortions is a core goal of cognitive behavioural therapy. Crucially, these distortions are primarily expressed through language, making them well-suited to computational modelling. Recent work in natural language processing has begun to explore the automatic detection and classification of cognitive distortions, with applications ranging from clinical decision support tools to mental health chatbots, journaling tools, and triage systems. Studies have shown that incorporating cognitive distortion-level features can improve outcomes in related tasks such as depression detection (Wang et al., 2023b), complementing more traditional sentiment or topic based approaches. By identifying distorted cognitive patterns in everyday text, natural language processing systems may support more timely, personalised, and psychologically-informed interventions.
@@
To our knowledge, the most comparable prior survey is by Suputra et al. (2023), which examined 12 studies and provided an initial synthesis of modelling approaches used at the time. Our work expands on this by covering 38 publications, including recent preprints, and offering a structured overview of the computational landscape for cognitive distortion detection and classification. This survey aims to provide a clear and practical entry point for researchers engaged in the growing field of cognitive distortion detection and classification from a computational perspective. Our focus is firmly on how computational methods approach these tasks, without seeking to redefine or adjudicate psychological constructs themselves. The contributions in this paper are threefold. (i) It provides a consolidated reference of the cognitive distortion taxonomies used across computational studies, highlighting inconsistencies and listing common synonyms (Tables 1 and 3). (ii) It defines and analyses task setups, datasets, computational methods, and approaches to performance evaluation, highlighting key patterns and gaps (§2, §4, §5, §6). (iii) It identifies open challenges (§7) and proposes best practices to guide future research (§8). In doing so, this paper aims to enable more consistent, comparable, and reproducible work in this emerging field.
@@ 3 Taxonomies of cognitive distortions in natural language processing
Despite the shared objective of classifying cognitive distortions, studies seeking to label cognitive distortions have adopted diverse taxonomies. Without standardisation, it is hard to keep annotations consistent, compare models properly, or clearly interpret the results. Table 1 illustrates this fragmentation by mapping which cognitive distortion categories are recognised across the papers surveyed. While some categories, such as Labelling and Should Statements, are commonly used, others are inconsistently applied or redefined in various ways. Early conceptions of cognitive distortions were proposed by Beck (1963), who described patterns of dysfunctional thinking with examples such as arbitrary inference and overgeneralisation. This work was later popularised by Burns (1999), whose ten-category taxonomy is frequently cited in psychology literature. [^3] While Burns (1999) is not the only taxonomy of cognitive distortions used in psychological contexts, and clinical practice often employs broader or alternative frameworks, our focus is on the inconsistency in how these definitions are interpreted and applied within computational research. However, computational studies do not uniformly follow this framework.
@@ 4.3 Annotation Strategies
Annotation strategies vary considerably across datasets. Clinical and literature-derived corpora typically rely on expert annotators, prioritising label quality at the expense of scalability. In contrast, social media and crowd-sourced datasets often involve non-expert annotators, sometimes supported by brief, domain-specific training from qualified psychologists to improve consistency. While these strategies enable large-scale annotation, reliably classifying cognitive distortions remains challenging. The task demands subtle, often subjective judgements, and studies consistently report low inter-annotator agreement, particularly when annotators lack deeper domain expertise. To mitigate this, some works adopt strict inclusion criteria, only retaining examples where annotators fully agree on the label or a subset of distortion types (Aureus et al., 2021; Shickel et al., 2020). Though this approach improves label precision, it risks introducing bias by systematically excluding ambiguous or borderline cases - which are arguably the most reflective of real-world cognitive distortion occurrences.
@@ 5 Modelling Approaches
This section outlines computational approaches to cognitive distortion detection and classification, grouped into six methodological categories that reflect major developments in the field.
@@ 5.2 Traditional Machine Learning (Feature-based)
As more annotated corpora became available, early rule-based systems gave way to feature-based statistical models. These approaches combined classic classifiers, such as logistic regression (LR) and support vector machines (SVMs), with engineered features such as Linguistic Inquiry and Word Count (LIWC) scores and Term Frequency-Inverse Document Frequency (TF-IDF) vectors. Simms et al. (2017) demonstrated that LR trained on LIWC features performed well on the detection task using Tumblr data, while Shickel et al. (2020) found that TF-IDF with LR outperformed CNNs on a 15-class single-label synthetic dataset (F1 = 0.68), indicating that shallow linguistic cues can remain competitive even in classification tasks. Similarly, Shreevastava and Foltz (2021) used SVM with smooth inverse frequency (SIF) embeddings, achieving strong performance (F1 = 0.77) on the detection task despite the insensitivity of SIF to word order.
@@ 5.3 Deep Learning with Static Embeddings
Static word embeddings such as Word2Vec (Mikolov et al., 2013) and GloVe (Pennington et al., 2014) introduced vector-based representations of words derived from co-occurrence patterns. These effectively captured word similarity but failed to account for contextual nuance (e.g., ‘riverbank’ vs ‘bank’). Despite this limitation, such embeddings formed the backbone of early deep-learning based cognitive distortion classifiers. For example, Rojas-Barahona et al. (2018) combined GloVe with CNNs for multi-label cognitive distortion classification, outperforming traditional models. Similarly, Mostafa et al. (2021) trained LSTM models using GloVe vectors to classify different cognitive distortion types, although their findings highlighted concerns around overfitting due to the use of limited and synthetic data. To address data sparsity in Arabic texts, Alhaj et al. (2022) applied contextual topic modelling (CTM), integrating static embeddings with domain-specific topic information via BERTopic - an approach that proved helpful in low-resource settings. While static embeddings offered richer representations than earlier methods, their inability to capture context, especially for polysemous words and subtle pragmatic distinctions, ultimately led to the rise of contextual models.
@@ 5.4 Transformer-based Architectures
Transformers, particularly BERT (Bidirectional Encoder Representations from Transformers) (Devlin et al., 2019), introduced contextual embeddings that capture word meaning based on the surrounding sentence, marking a significant leap forward in many natural language processing tasks, including cognitive distortion detection and classification. Shreevastava and Foltz (2021) used fine-tuned Sentence-BERT (SBERT) for binary cognitive distortion detection, showing notable improvements over earlier models. Similarly, Lybarger et al. (2022) demonstrated that incorporating conversational history further improves classification performance on therapy dialogues. Domain-adapted transformers like MentalBERT (Ji et al., 2022), which are pre-trained on mental-health-related data, showed benefits over general-purpose models, while augmentation techniques such as mixup (Zhang et al., 2018) were explored to improve performance on rare cognitive distortion classes (minor improvements). Maddela et al. (2023) found that models like RoBERTa performed better than larger language models such as GPT-3.5 for this task, though distinguishing between closely related distortions remained a challenge. Overall, transformers brought significant gains in robustness and accuracy, but these came at the cost of reduced interpretability and a higher risk of overfitting, especially in datasets with extreme class imbalance. These limitations set the stage for the emergence of prompt-based models.
@@ 5.6 Multimodal and Multi-task Architectures
5.7 Feasibility of Meta-Analysis
While we initially intended to include comparative performance tables, meaningful aggregation proved infeasible. The studies diverge across critical axes - (i) task formulation, (ii) cognitive distortion taxonomy choice, (iii) dataset domain, (iv) evaluation metrics, (v) granularity of the unit of analysis, (vi) context inclusion and window size, and (vii) modality - often incompatibly. Sample sizes within aligned subgroups are too small for robust comparison, and inter-annotator agreement is reported inconsistently, further limiting comparability. Pooled tables would therefore risk suggesting misleading trends. We release the full set of extracted results in our GitHub repository, enabling researchers to build their own comparisons. In this paper, we restrict ourselves to qualitative synthesis, deferring formal meta-analysis until the evidence base is larger and more standardised.
@@ 6 Evaluation
Despite progress in cognitive distortion classification, evaluation practices remain inconsistent, with studies differing markedly in their choice and reporting of metrics, which hinders comparability. Although F1 score is the most commonly used metric, distinctions between macro, micro, and weighted variants are frequently overlooked - a significant issue for class-imbalanced datasets where per-class performance is critical. Some studies now report AUPRC (Area Under the Precision-Recall Curve) to account for skewed label distributions, providing a more informative measure of performance on rare cognitive distortion types (Ding et al., 2022). Nonetheless, per-class metrics are still underreported, hiding weaknesses in addressing infrequent distortions. Similarly, dataset quality is often inconsistently assessed. Inter-annotator agreement is either inconsistently measured or reported using incomparable metrics. While Cohen’s Kappa ($\kappa$), which adjusts for chance, is typically more appropriate, many studies instead rely on raw agreement or non-standard metrics, blurring the line between annotation reliability and model performance. Baseline comparisons are further complicated by inconsistent cognitive distortion taxonomies and datasets.
@@ 7.1 Inconsistent cognitive distortion Taxonomies
A longstanding challenge is the lack of a standardised taxonomy for cognitive distortions. While foundational frameworks such as the Burns ten-category list (Burns, 1999) are commonly cited, computational studies diverge significantly in how they define, split, or rename distortion types. For instance, Jumping to Conclusions is frequently subdivided into Mind Reading and Fortune Telling, and terms such as All or Nothing Thinking appear under multiple aliases (e.g., Black and White Thinking, Polarised Thinking), making it difficult to compare models across studies, reproduce results, or interpret outputs reliably. These inconsistencies also affect annotation quality, as ambiguous or overly granular label sets introduce subjectivity and reduce inter-annotator agreement - an issue compounded by the lack of formal guidance on taxonomy use.
@@ 7.2 Data Scarcity, Imbalance, and Annotation Limits
The field remains constrained by a lack of large, high-quality datasets that capture authentic, context-rich examples of distorted thinking. Clinical corpora are scarce and often inaccessible due to privacy constraints, while many widely used datasets are synthetic, crowd-sourced, or compiled from multiple domains, which may lack the nuance and ambiguity of real-world language. This limits the complexity of distortions that models can learn and typically results in heavy class imbalance, with rare distortion types being underrepresented or excluded altogether. Although augmentation strategies such as mixup or back-translation offer minor gains for rare categories (Ding et al., 2022), a deeper issue lies in the ceiling imposed by annotation reliability itself. In some settings, even expert annotators show limited agreement, particularly for subtle or overlapping categories. For instance, Tauscher et al. (2023) report that for the presence of ‘Any Distortion’ in a text, human F1 agreement was 0.63, while a fine-tuned BERT model reached 0.62 - a very small difference. This suggests that for certain formulations, such as binary detection or high-frequency classes, current models may already be approaching the upper bound set by annotation quality. It also reinforces the need for clearer task definitions and more consistent annotation protocols before investing in model complexity.
@@ 7.3 Overreliance on Short Text
The vast majority of existing datasets frame cognitive distortion detection at the sentence or single-utterance level. This simplifies annotation and model design but introduces strong limitations, as many distortions are context-dependent or only weakly signalled lexically. By stripping away discourse-level cues, models are forced to rely on surface-level patterns and may perform poorly on more ambiguous cases. Predictably, this has contributed to the dominance of distortion types with overt markers (e.g., Should Statements) in both datasets and model outputs. Empirical studies confirm the value of richer context, showing that including prior conversational turns improved detection F1 from 0.68 to 0.73 (Lybarger et al., 2022), while frameworks such as ERD achieve greater interpretability by explicitly reasoning over multi-sentence inputs (Lim et al., 2024). Still, most benchmarks continue to prioritise short-form inputs. Moving forward, we argue that context-aware models should become the norm rather than the exception. In parallel, new datasets should prioritise multi-turn conversations, real patient narratives, and longer-form content that more closely mirrors therapeutic language.
8 Best Practices and Recommendations
Alongside consolidating existing research, it is important to address the main sources of fragmentation in the field. We therefore propose a set of best practices for future work, which should be followed where possible or clearly justified if deviated from.
8.1 Taxonomy Adoption
As shown in Table 1, inconsistent use of cognitive distortion taxonomies has made cross-study comparison difficult. In the absence of a universally accepted taxonomy, we recommend Burns’ taxonomy [^4] We provide recommended label sets online. (Burns, 1999) as a sensible default, since it is the most widely cited and most alternatives used in natural language processing are partial reinterpretations of it. Researchers should (i) report the source and rationale for their chosen taxonomy, (ii) avoid introducing new or expanded taxonomies without justification, and (iii) prioritise taxonomies grounded in clinical consensus. Reliance on loosely defined online taxonomies is discouraged. Where deviations from Burns’ taxonomy are necessary, the rationale should be documented in the study, or, in the case of new datasets, in the corresponding datasheet, for which we provide a template online.
8.2 Unambiguous Evaluation Reporting
Inconsistent reporting is a major barrier to comparing results across studies. To improve comparability, we recommend that future work (i) clearly state the task formulation (detection, single-label, or multi-label classification; §2), (ii) specify the analysis unit (sentence, turn, session) and, if relevant, document the exact context window, (iii) indicate the cognitive distortion taxonomy used, (iv) report per-class scores alongside macro and weighted F1555For imbalanced or multi-label settings, AUPRC (micro or macro) and per-class PR curves may be more informative. Ultimately, metrics should fit the use case. with unambiguous labels, and (v) explain the choice of evaluation metrics. While a single metric cannot suit all applications, departures from macro or weighted F1 should always be accompanied by a clear rationale.
8.3 Dataset Development and Use
For researchers creating new cognitive distortion datasets, we recommend providing a datasheet that documents the dataset’s origin, annotation protocol, size, taxonomy, analysis unit, and licensing. To support this, we provide a standardised datasheet template in our GitHub repository. For reuse of existing datasets, we encourage researchers to apply our Ethics Flowchart, which provides practical guidance on assessing provenance, consent, and documentation before experimentation.
8.4 Annotation Reliability and Inter-annotator Agreement
Annotation processes should be reported transparently. Researchers should provide standard inter-annotator agreement metrics (e.g., Cohen’s $\kappa$, Fleiss’ $\kappa$, Krippendorff’s $\alpha$) rather than vague statements of ‘agreement’. The rationale for the chosen metric should be stated, and partial dataset sampling for inter-annotator agreement is acceptable provided that procedures are clearly documented.
We further recommend that future work (i) reports human-model performance comparison metrics where possible, (ii) prioritises the release of multilingual and multi-domain corpora, and (iii) explicitly documents how annotation disagreements are resolved. Without such practices, gains in model performance may reflect noise-fitting rather than genuine progress.
8.5 Code and Dataset Release
To support replication and benchmarking, researchers should release code and, where licensing and privacy considerations permit, datasets. As shown in Table 4, many of the surveyed studies do not provide public implementations. We recommend that future work adopt code and data release as standard practice, in line with recent broader calls for stronger reproducibility standards in AI research and governance (Semmelrock et al., 2025; Mason-Williams and Mason-Williams, 2025).